OpenStudy (anonymous):

Radio signals travel at a rate of 3 x 10^8 meters per second. How many seconds would it take for a radio signal to travel from a satellite to the surface of Earth if the satellite is orbiting at a height of 9.6 x 10^6 meters? (Hint: Time is distance divided by speed.)

OpenStudy (anonymous):

A. 3.2 x 10^2 seconds B. 3.2 x 10^-2 seconds C. 3.13 x 10^1 seconds D. 2.88 x 10^15

OpenStudy (anonymous):

*seconds. I meant to add seconds to letter D :P

OpenStudy (abhisar):

\(\sf \boxed{\huge Time=\frac{Distance}{Speed}}\)

OpenStudy (anonymous):

Okay, that helps me a little bit...

OpenStudy (abhisar):

Distance is given \(\sf 9.6 \times 10^6 m\) and speed of signal is given \(\sf 3.8 \times 10^8 m/s\). Now all you have to do is plug in the values and find the answer.